OpenStudy (anonymous):

An object weighs 39.36 newtons. What is its mass if a gravitometer indicates that g = 9.83 m/sec2?

OpenStudy (anonymous):

F = mg, force is 39.96 newtons, g is 9.83, so mass = 9.83/39.96

OpenStudy (anonymous):

\[Wg = mass * gravity\] \[mass = \frac{ Wg }{ gravity } = \frac{ 39.36 }{ 9.83 } = ~4Kg\]